# Auto-GDA: Automatic Domain Adaptation For Grounding Verification in Retrieval Augmented Generation

### v1

@@ Abstract
While retrieval augmented generation (RAG) has been shown to enhance factuality of large language model (LLM) outputs, LLMs still suffer from hallucination, generating incorrect or irrelevant information. One common detection strategy involves prompting the LLM again to assess whether its response is grounded in the retrieved evidence, but this approach is costly. Alternatively, lightweight natural language inference (NLI) models for efficient grounding verification can be used at inference time. While existing pre-trained NLI models offer potential solutions, their performance remains subpar compared to larger models on realistic RAG inputs. RAG inputs are more complex than most datasets used for training NLI models and have characteristics specific to the underlying knowledge base, requiring adaptation of the NLI models to a specific target domain. Additionally, the lack of labeled instances in the target domain makes supervised domain adaptation, e.g., through fine-tuning, infeasible. To address these challenges, we introduce Automatic Generative Domain Adaptation (Auto-GDA). Our framework enables unsupervised domain adaptation through synthetic data generation. Unlike previous methods that rely on handcrafted filtering and augmentation strategies, Auto-GDA employs an iterative process to continuously improve the quality of generated samples using weak labels from less efficient teacher models and discrete optimization to select the most promising augmented samples. Experimental results demonstrate the effectiveness of our approach, with models fine-tuned on synthetic data using Auto-GDA often surpassing the performance of the teacher model and reaching the performance level of LLMs at 10 % of their computational cost.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly used in consequential applications. Despite their versatility, LLMs often produce hallucinations, in which the generated information is inaccurate or fabricated and require costly retraining to integrate new knowledge. One promising method to mitigate these issues is retrieval augmented generation (RAG, Lewis et al., 2020). RAG enhances text generation by adding information from external knowledge sources to the prompt and has been shown to reduce hallucinations in practice (Shuster et al., 2021). Nevertheless, even when the most capable LLMs are used with RAG, hallucination rates of 20 – 30% persist (Santhanam et al., 2021).
To prevent hallucinated output from being delivered to end-users, natural language inference (NLI) models can be used to verify grounding of the generated output in the documents retrieved (Chen et al., 2023b; Es et al., 2024; Tang et al., 2024) before the output is relayed to the end-user: the generated response must be fully grounded in the documents, i.e., it must be logically inferrable from the documents; otherwise, it is considered ungrounded. However, as we need
RAG task performance (ROC-AUC)inference time lightweight: e.g., DeBERTa, Mini-Check Pre/Post-Processing: e.g., AlignScore LLMs: e.g., GPT-4, Claude-3.5 domain-adapted, lightweight model 1x2.5x10x0.72absent0.72$\approx$ 0.72≈ 0.72 +20%
@@
to check the outputs at inference time, we require lightweight NLI models with very low latency. The current landscape of available NLI models for verifying grounding in RAG is illustrated in Figure 1 based on results obtained in our evaluation of correctness and inference time (see Table 3 for full numeric results): Some recent works such as Mini-Check (Tang et al., 2024) have developed lightweight models for NLI, e.g., based on RoBERTa (Liu, 2019). These models have shown good performance on academic benchmarks. However, our results indicate that their performance in verifying grounding for realistic RAG inputs lags behind LLMs by about 20% (in ROC-AUC scores). Other recent methods use pre- and post-processesing techniques such as sentence tokenization or LLM prompting to decompose long prompts (Zha et al., 2023; Es et al., 2024) into several chunks or facts. Each of these chunks needs to be processed in a separate forward pass, resulting in high latency as well. While some studies (e.g., Manakul et al., 2023; Tang et al., 2024) have also explored directly using LLMs like GPT-4 for text entailment detection, their latency is about an order of magnitude above the lightweight models. Taken together, these characteristics make it hard to deploy the existing approaches in real-time industry use-cases.
@@
Bridging this domain gap poses a significant challenge due to the inherent difficulty of adapting models to unseen domains that is further amplified by the prohibitive costs of obtaining labeled data from the target domain. This prevents supervised domain adaptation, e.g., through fine-tuning on target domain data. To address this issue, we propose Automatic Generative Domain Adaptation (Auto-GDA). Our unsupervised domain adaptation framework produces high-quality synthetic data, which is then used to fine-tune a lightweight NLI model, adapting it to a specific domain of RAG inputs. While training data generation by simply prompting LLMs has been repeatedly explored in the literature (e.g., Saad-Falcon et al. (2024); Hosseini et al. (2024)), data quality might be further improved through filtering and incorporating background knowledge through label-preserving data augmentation strategies, such as round-trip translation (Chen et al., 2023b). However, specifying good filters and heuristic augmentation strategies require significant manual effort. As data augmentations can further be applied iteratively, the space of potential samples grows exponentially, necessitating efficient search strategies. During this offline training phase, less efficient teacher models can provide additional guidance using weak labels. Auto-GDA offers a unified way to leverage all these available tools. We thus make the following contributions:
@@
    We show that our objective corresponds to an enhanced distribution matching objective but is highly efficient to optimize.
@@
    Our experiments on realistic RAG inputs highlight that our fine-tuned models using Auto-GDA (1) often outperform their weak teacher models (2) perform almost as well as reference models fine-tuned with human-labeled data and (3) reach the level of performance exhibited by LLMs while having almost 10x lower latency. (4) Our method further outperforms more classical training-based unsupervised domain adaptation techniques.
@@ 2 Related Work
Synthetic Data for Domain Adaptation in NLP. While synthetic QA data generation is well-explored (Shakeri et al., 2020; Ushio et al., 2022; Yue et al., 2022; Lee et al., 2023), synthetic data for NLI domain adaptation has received less attention, potentially due to the difficulty of generating realistic and difficult samples. Wang et al. (2023) propose an iterative synthetic data generation scheme requiring partially labeled data. They generate initial seed data using an LLM prompt that is iteratively refined based on errors from a model trained on human-labeled reference set. Unlike this work, we assume very limited access to labeled data from the target domain.
Knowledge Distillation. We borrow the term “teacher model” from the knowledge distillation literature (Gou et al., 2021; Yang et al., 2020). However, our problem differs from distillation problems because our target dataset is unlabeled.
In this paper, we focus on the problem of systematically generating and selecting the most beneficial synthetic samples that can be created through initial generation and iterative augmentation steps. We do so using an efficient objective that can be interpreted as a form of distribution matching.
@@ 3 Preliminaries
Domain adaptation is concerned with adapting an ML model pretrained on a source domain to make predictions on a target domain when the underlying data distributions differ across the two domains. The unsupervised domain adaptation problem is further complicated due to the lack of labeled data in the target domain. This means that while features are available, there is no direct information about the correct class labels for the target domain samples. This poses a significant challenge as the model must learn to adapt to the new distribution without explicit guidance.
@@ 3.1 Unsupervised Domain Adaptation for NLI
In this work, we focus on covariate shift between the two domains: While the prior $p({\bm{e}},{\bm{c}})$ is subject to change across domains, the true relation between specific features and labels, $p(y|{\bm{e}},{\bm{c}})$ is consistent for the source and the target domain. For the NLI task considered here, this assumption is sensible because the entailment relation itself does not change for different domains. Following prior work Saad-Falcon et al. (2024), we slightly deviate from the fully unsupervised setup by supposing that a very small portion of the target domain can be manually labeled and used as a validation set for hyperparameter tuning only, as is commonly done in NLI literature (Laban et al., 2022; Tang et al., 2022; Zha et al., 2023). We show that our method works with validation sets as small as $30$ samples.
@@ 4.1 Outline of the framework
Auto-GDA is specifically designed to tackle these three challenges. As RAG outputs stem from LLMs, we also generate realistic initial samples using LLMs. We leverage few-shot prompting to transfer patterns in the output to the generated samples. To preserve the diversity of the evidence (which contains the relevant documents in the knowledge base), we generate synthetic claims sequentially for each unique piece of evidence ${\bm{e}}$ available in the unlabeled target dataset $\mathcal{D}_{t}$. This has the advantage that a broad diversity of documents in the knowledge base is represented. We propose to apply augmentations on the synthetic data to increase diversity further. As the augmentation strategies are only approximately label-preserving, we have to keep track of increasing label uncertainty to detect samples with low-quality labels when several data augmentation steps are applied. We therefore equip each sample with an entailment certainty score $r$, an estimate of the probability of the sample having an entailed label ($y{=}1)$ which can be used to remove samples with low-quality labels. Auto-GDA applies these steps iteratively to successively increase data quality. In summary our framework consists of the following steps, which we describe in more detail in the next sections:
@@
    Initial Generation. Generate an initial sample population $\mathcal{D}^{(0)}_{\bm{e}}=\{(\hat{{\bm{c}}}_{k},\hat{y}_{k},r_{k}^{(0)})\}_{k=1}^{K}$ of claims $\hat{{\bm{c}}}$ and labels $\hat{y}$ for the evidence ${\bm{e}}$ using the generator $G$. Use the teacher model $T$ to assign initial entailment certainty $r^{(0)}$ scores to each sample of synthetic data. This results in each sample having a hard label $\hat{y}$ and a “soft” confindence score $r^{(0)}$ for the hard label.
@@ 4.2 Generating Realistic Initial Data
LLMs have been repeatedly used to generate synthetic data for various domains, including NLI (Saad-Falcon et al., 2024; Hosseini et al., 2024). In this work, we generate initial data using few-shot prompting with the prompts provided in Section E.1. The prompt instructs the LLM to generate synthetic claims $\hat{{\bm{c}}}=G({\bm{e}},\text{claim}(\mathcal{D}_{t,{\bm{e}}}),\hat{y})$ for the evidence ${\bm{e}}$, reflecting the style of example claims from $\mathcal{D}_{t}$ ($\text{claim}(\mathcal{D}_{t,{\bm{e}}})$ denoting claims from target data for the evidence ${\bm{e}}$) and target label $\hat{y}\in 0,1$. For label $\hat{y}=1$, the LLM is instructed to include only grounded facts, for $\hat{y}=0$, some ungrounded information should be introduced. We assign labels $\hat{y}$ according to the prompt used, resulting in complete initial generated tuples $(\hat{{\bm{c}}},\hat{y})$. We follow some related works (Puri et al., 2020; Vu et al., 2021), which have suggested generating many samples and only keeping the most confident. To do so, the samples can be equipped with a weak estimate of the label probability using the teacher model, e.g., another LLM or an NLI model with sophisticated pre- and postprocessing. In the binary classification setup, we can compute initial entailment certainties as $r^{(0)}=T({\bm{e}},\hat{{\bm{c}}})$, which can be interpreded as an uncalibrated and potentially noisy estimate of $p(y=1|{\bm{e}},\hat{{\bm{c}}})$. We explore LLMs for data generation and use state-of-the-art NLI models and also LLMs as teacher models $T$ for providing initial entailment certainties. Adding the entailment certainty scores $r^{(0)}$ to the respective tuples we obtain a set of triples $\mathcal{D}^{(0)}_{\bm{e}}=\{(\hat{{\bm{c}}}_{k},\hat{y}_{k},r_{k}^{(0)})\}_{k=1}^{K}$ after this step.
@@ 4.3 Increasing Diversity through Label-Preserving Data Augmentations
using the entailment certainty $r^{(i)}$ of the original tuple $({\bm{e}},\hat{{\bm{c}}})$ as a base. Note that some teacher models may be particularly reliable with claim-claim pairs than with evidence-claim pairs so it can be useful to choose a different teacher model for this update than for computing initial certainty scores.
Label Invariant Augmentation Strategies: In this work, we consider three augmentation strategies that will likely preserve entailment labels (see Appendix Section C.1 for additional details):
  • Partial Rephrasing with LLMs. Our first augmentation is an LLM-based rephrasing step. Specifically, we randomly mask 20% of the words of the input sequence by replacing the corresponding words by “_" and ask an LLM (Claude3 Haiku) to impute the gaps while preserving the meaning.
  • Complete Paraphrasing. We use a T5-based paraphrasing model Vorobev & Kuznetsov (2023). We generate paraphrases for the claims using enforcing diversity using a constraint that prevents $n$-grams of length greater than 5 from being regenerated.

### v2

# Auto-GDA: Automatic Domain Adaptation for Grounding Verification in Retrieval-augmented Generation

@@ Abstract
While retrieval-augmented generation (RAG) has been shown to enhance factuality of large language model (LLM) outputs, LLMs still suffer from hallucination, generating incorrect or irrelevant information. A common detection strategy involves prompting the LLM again to assess whether its response is grounded in the retrieved evidence, but this approach is costly. Alternatively, lightweight natural language inference (NLI) models for efficient grounding verification can be used at inference time. While existing pre-trained NLI models offer potential solutions, their performance remains subpar compared to larger models on realistic RAG inputs. RAG inputs are more complex than most datasets used for training NLI models and have characteristics specific to the underlying knowledge base, requiring adaptation of the NLI models to a specific target domain. Additionally, the lack of labeled instances in the target domain makes supervised domain adaptation, e.g., through fine-tuning, infeasible. To address these challenges, we introduce Automatic Generative Domain Adaptation (Auto-GDA). Our framework enables unsupervised domain adaptation through synthetic data generation. Unlike previous methods that rely on handcrafted filtering and augmentation strategies, Auto-GDA employs an iterative process to continuously improve the quality of generated samples using weak labels from less efficient teacher models and discrete optimization to select the most promising augmented samples. Experimental results demonstrate the effectiveness of our approach, with models fine-tuned on synthetic data using Auto-GDA surpassing the performance of the teacher model and reaching the performance level of LLMs at 10 % of their computational cost.
**footnotetext: Work done during internship at AWS AI Labs.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly used in consequential applications. Despite their versatility, LLMs often produce hallucinations, in which the generated information is inaccurate or fabricated and require costly retraining to integrate new knowledge. One promising method to mitigate these problems is retrieval-augmented generation (RAG, Lewis et al., 2020). RAG enhances text generation by adding information from external knowledge sources to the prompt and has been shown to reduce hallucinations in practice (Shuster et al., 2021). Nevertheless, even when modern LLMs are used with RAG, hallucination rates of 15% – 30% (Chen et al., 2023a) or more than one hallucination per 100 output tokens can occur (Niu et al., 2024).
To prevent hallucinated output from being delivered to end-users, natural language inference (NLI) models can be used to verify the grounding of the generated output in the documents retrieved (Chen et al., 2023b; Es et al., 2024; Tang et al., 2024) before the output is relayed to the end-user: the
grounding verification performance (RAG data, ROC-AUC)inference time lightweight: e.g., DeBERTa, Mini-Check Pre/Post-Processing: e.g., AlignScore LLMs: e.g., GPT-4, Claude-3.5 domain-adapted, lightweight model 1x2.5x10x0.72absent0.72$\approx$ 0.72≈ 0.72 +20%
@@
generated response must be fully grounded in the documents, i.e., it must be logically inferrable from the documents; otherwise, it is considered ungrounded. However, as we need to check the outputs at inference time, we require lightweight NLI models with very low latency. The current landscape of available NLI models for verifying grounding in RAG is illustrated in Figure 1 based on results obtained in our evaluation of correctness and inference time (see Table 3 for full numeric results): Some recent works such as Mini-Check (Tang et al., 2024) have developed lightweight models for NLI, e.g., based on RoBERTa (Liu, 2019). These models have shown good performance on academic benchmarks. However, our results indicate that their performance in verifying grounding for realistic RAG inputs lags behind LLMs by about 20% (in ROC-AUC scores). Other recent methods use pre- and post-processesing techniques such as sentence tokenization or LLM prompting to decompose long prompts (Zha et al., 2023; Es et al., 2024) into several chunks or facts. Each of these chunks needs to be processed in a separate forward pass, resulting in high latency as well. While some studies (e.g., Manakul et al., 2023; Tang et al., 2024) have also explored directly using LLMs like GPT-4 for text entailment detection, their latency is about an order of magnitude above the lightweight models. Taken together, these characteristics make it hard to deploy the existing approaches in real-time industry use-cases.
@@
Bridging this domain gap poses a significant challenge due to the inherent difficulty of adapting models to unseen domains that is further amplified by the prohibitive costs of obtaining labeled data from the target domain. This prevents supervised domain adaptation, e.g., through fine-tuning on target domain data. To address this issue, we propose Automatic Generative Domain Adaptation (Auto-GDA). Our unsupervised domain adaptation framework produces high-quality synthetic data, which is then used to fine-tune a lightweight NLI model, adapting it to a specific domain of RAG inputs. While training data generation by simply prompting LLMs has been repeatedly explored in the literature (e.g., Saad-Falcon et al., 2024; Hosseini et al., 2024), data quality might be further improved through filtering and incorporating background knowledge through label-preserving data augmentation strategies, such as round-trip translation (Chen et al., 2023b). However, specifying good filters and heuristic augmentation strategies require significant manual effort. As data augmentations can further be applied iteratively, the space of potential samples grows exponentially, necessitating efficient search strategies. During this offline training phase, less efficient teacher models can provide additional guidance using weak labels. Auto-GDA offers a unified way to leverage all these available tools. We thus make the following contributions:
@@
    We show that our objective corresponds to an enhanced distribution matching objective but is highly efficient to optimize. w
@@
    Our experiments on realistic RAG inputs highlight that our fine-tuned models using Auto-GDA (1) often outperform their weak teacher models (2) perform almost as well as reference models fine-tuned with human-labeled data and (3) reach the level of performance exhibited by LLMs while having almost 10x lower latency. (4) Our method further outperforms more classical training-based unsupervised domain adaptation techniques. [^1] Code is available at https://github.com/amazon-science/AutoGDA-Efficient-Grounding-Verification-in-RAG
@@ 2 Related Work
Synthetic Data for Domain Adaptation in NLP. While synthetic QA data generation is well-explored (Shakeri et al., 2020; Ushio et al., 2022; Yue et al., 2022; Lee et al., 2023), synthetic data for NLI domain adaptation has received less attention, potentially due to the difficulty of generating realistic and difficult samples. Wang et al. (2023a) propose an iterative synthetic data generation scheme requiring partially labeled data. They generate initial seed data using an LLM prompt that is iteratively refined based on errors from a model trained on human-labeled reference set. Unlike this work, we assume very limited access to labeled data from the target domain.
Classical Unsupervised Domain Adaptation. Beyond synthetic data approaches, classical unsupervised domain adaptation (UDA) techniques have also been applied in NLP. These include techniques to align embeddings (Chen et al., 2018; Li et al., 2018; Choudhry et al., 2022; Sun & Saenko, 2016), pre-training approaches (Gururangan et al., 2020; Han & Eisenstein, 2019) or virtual adversarial training (Miyato et al., 2019; Jiang et al., 2020). See Appendix A for details.
We borrow the term “teacher model” from the knowledge distillation literature (Gou et al., 2021; Yang et al., 2020). However, our problem differs from distillation problems because our target dataset is unlabeled. In this paper, we focus on the problem of systematically generating and selecting the most beneficial synthetic samples that can be created through initial generation and iterative augmentation steps. We do so using an efficient objective that can be interpreted as a form of distribution matching.
@@ 3 Preliminaries
Domain adaptation is concerned with adapting an ML model pretrained on a source domain to make predictions on a target domain when the underlying data distributions differ across the two domains. The unsupervised domain adaptation problem is further complicated due to the lack of labeled data in the target domain. While features are available, there is no direct information about the correct class labels for the target domain samples. This poses a significant challenge as the model must learn to adapt to the new distribution without explicit guidance.
@@ 3.1 Unsupervised Domain Adaptation for NLI
In this work, we focus on covariate shift between the two domains: While the prior $p({\bm{e}},{\bm{c}})$ is subject to change across domains, the true relation between specific features and labels, $p(y|{\bm{e}},{\bm{c}})$ is consistent for the source and the target domain. For the NLI task considered here, this assumption is sensible because the entailment relation itself does not change for different domains. Following prior work (Saad-Falcon et al., 2024), we slightly deviate from the fully unsupervised setup by supposing that a very small portion of the target domain can be manually labeled and used as a validation set for hyperparameter tuning only, as is commonly done in NLI literature (Laban et al., 2022; Tang et al., 2022; Zha et al., 2023). We show that our method works with validation sets as small as $30$ samples.
@@ 4.1 Outline of the framework
Auto-GDA is specifically designed to tackle these three challenges. As RAG outputs stem from LLMs, we also generate realistic initial claim samples for a given evidence using LLMs. We leverage few-shot prompting to transfer patterns in the output to the generated samples. To preserve the diversity of the evidence (which contains the relevant documents in the knowledge base), we generate synthetic claims sequentially for each unique piece of evidence ${\bm{e}}$ available in the unlabeled target dataset $\mathcal{D}_{t}$. This has the advantage that the broad diversity of documents in the knowledge base is represented. We propose to apply augmentations on the synthetic data to increase diversity further. As the augmentation strategies are only approximately label-preserving, we have to keep track of increasing label uncertainty to detect samples with low-quality labels when several data augmentation steps are applied. We therefore equip each sample with an entailment certainty score $r$, an estimate of the probability of the sample having an entailed label ($y{=}1)$ which can be used to remove samples with low-quality labels. Auto-GDA applies these steps iteratively to successively increase data quality. In summary our framework consists of the following steps, which we describe in more detail in the next sections:
@@
    Initial Generation. Generate an initial sample population $\mathcal{D}^{(0)}_{\bm{e}}=\{(\hat{{\bm{c}}}_{k},\hat{y}_{k},r_{k}^{(0)})\}_{k=1}^{K}$ of claims $\hat{{\bm{c}}}$ and labels $\hat{y}$ for the evidence ${\bm{e}}$ using the generator $G$. Use the teacher model $T$ to assign initial entailment certainty $r^{(0)}$ scores to each sample of synthetic data. This results in each sample having a hard label $\hat{y}$ and a “soft” confidence score $r^{(0)}$ for the hard label.
@@ 4.2 Generating Realistic Initial Data
LLMs have been repeatedly used to generate synthetic data for various domains, including NLI (Saad-Falcon et al., 2024; Hosseini et al., 2024). In this work, we generate initial data using few-shot prompting with the prompts provided in Section E.1. The prompt instructs the LLM to generate synthetic claims $\hat{{\bm{c}}}=G({\bm{e}},\text{claim}(\mathcal{D}_{t,{\bm{e}}}),\hat{y})$ for the evidence ${\bm{e}}$, reflecting the style of example claims from $\mathcal{D}_{t}$ ($\text{claim}(\mathcal{D}_{t,{\bm{e}}})$ denoting claims from target data for the evidence ${\bm{e}}$) and target label ${\hat{y}\in\{0,1\}}$. For label $\hat{y}=1$, the LLM is instructed to include only grounded facts, for $\hat{y}=0$, some ungrounded information should be introduced. We assign labels $\hat{y}$ according to the prompt used, resulting in complete initial generated tuples $(\hat{{\bm{c}}},\hat{y})$. We follow some related works (Puri et al., 2020; Vu et al., 2021), which have suggested generating many samples and only keeping the most confident. To do so, the samples can be equipped with a weak estimate of the label probability using the teacher model, e.g., another LLM or an NLI model with sophisticated pre- and postprocessing. In the binary classification setup, we can compute initial entailment certainties as $r^{(0)}=T({\bm{e}},\hat{{\bm{c}}})$, which can be interpreded as an uncalibrated and potentially noisy estimate of $p(y=1|{\bm{e}},\hat{{\bm{c}}})$. We explore LLMs for data generation and use state-of-the-art NLI models and also LLMs as teacher models $T$ for providing initial entailment certainties. Adding the entailment certainty scores $r^{(0)}$ to the respective tuples we obtain a set of triples $\mathcal{D}^{(0)}_{\bm{e}}=\{(\hat{{\bm{c}}}_{k},\hat{y}_{k},r_{k}^{(0)})\}_{k=1}^{K}$ after this step.
@@ 4.3 Increasing Diversity through Label-Preserving Data Augmentations
In this section, we demonstrate how to augment the initial synthetic dataset (generated using the few-shot prompting strategy) for additional diversity, while maintaining a high degree of label certainty for the augmented synthetic data points. We exploit a certain degree of background knowledge to derive data augmentation strategies (Chen et al., 2023b). For instance, we know that paraphrasing the claims while preserving their semantic meaning should not change their entailment label. However, when iteratively applying paraphrasing operations, we have to account for an increasing probability of accidentally flipping the label.
@@
using the entailment certainty $r^{(i)}$ of the original tuple $({\bm{e}},\hat{{\bm{c}}})$ as a base. Some teacher models may be particularly reliable with claim-claim pairs rather than with evidence-claim pairs, so it can be useful to choose a different teacher model for this update than for obtaining initial certainty scores.
Label Invariant Augmentation Strategies: In this work, we consider three augmentation strategies that will likely preserve entailment labels (see Section C.1 for additional details):
  • Partial Rephrasing with LLMs. Our first augmentation is an LLM-based rephrasing step. Specifically, we randomly mask 20% of the words of the input sequence by replacing the corresponding words by “_" and ask an LLM (Anthropic’s Claude3 Haiku) to impute the gaps while preserving the meaning.
  • Complete Paraphrasing. We use a T5-based paraphrasing model (Vorobev & Kuznetsov, 2023). We generate paraphrases for the claims enforcing diversity using a constraint that prevents $n$-grams of length greater than 5 from being regenerated.